# Symmetry-breaking induced surface magnetization in non-magnetic RuO2

### v1

@@ III.2.1 The emergence of surface magnetization in RuO2 (110) surface
Figure 2: Geometrical structure of the (110) surface and the emergence of surface magnetism at the surface. (a) The slab structure of 13L RuO2 (110) includes two Ru sublattices inherited from the bulk: the Ru-center sublattice (red), the Ru-corner sublattice (blue), and oxygen ligands represented in silver. At the surface layers, Ru-center atoms are fully octahedrally coordinated by six O ligands (denoted as Ru-6f), while the Ru-corner sublattice has one dangling bond due to surface termination (Ru-5f), see (b) for a better view. Also shown is the magnetization density visualization, which exhibits significant accumulation of spin moments only at the outermost layers, with spin-up in yellow and spin-down in cyan. The isosurfaces are plotted using 5% of the maximum isosurface value. (b) Detailed atomic structure of the surface includes stoichiometric surface termination, with labels for atoms in the surface layer. (c) The spin moment at each Ru site in each layer (top two panels), and the magnetization of each layer (bottom panel) are plotted as functions of distance from the bulk-like region. (d) A view along the [001] direction shows the layers numbered for the layer-projected DOS in (e), which demonstrates that spin polarization decays quickly when approaching the inner layers.
@@
To gain insights into the magnetic structure of the RuO2 (110) surface, we performed calculations using the slab model approach. In isostructural insulating materials such as TiO2, the (110) surface is electrostatically stable. However, electrostatics are not a concern for RuO2 since it is a metallic system. To simulate a surface with a bulk-like region well represented by a Ru layer in the middle of the structure, we chose stoichiometric slabs with an odd number of Ru layers. These slabs possess inversion symmetry, eliminating unnecessary complications related to asymmetric systems, such as spurious electric fields across the slab thickness. The slab thickness ranges from three layers (denoted as 3L) up to nineteen layers (19L). Each layer is composed of three atomic planes, with one Ru-containing layer sandwiched between two oxygen-only layers. Each slab has two identical surface layers at the top and bottom of the unit cell [cf. Fig. 2(a)].
@@
The converged, optimized thirteen-layer slab unit cell (13L) is shown in Fig. 2(a). This symmetric structure represents both bulk and surface areas of a semi-infinite bulk, as it is thick enough to decouple the two surfaces, and the middle layer behaves as a bulk-like region [see Fig. S1 in the Supplemental Material (SM) for the dependence of local moments of surface Ru atoms and top-layer magnetization on slab thickness]. Each Ru layer stacked along the [110] direction contains two Ru atoms corresponding to the Ru-center and Ru-corner sublattices of the bulk [cf. Fig. 1(a)]. This arrangement would yield a magnetically compensated structure with oppositely pointed moments within each (110) plane if the ground state of the bulk was antiferromagnetically ordered with the Néel vector along [001], as assumed earlier—similar to the case of the isostructural FeF2 antiferromagnet [50].
@@
Unlike the magnetic structure in traditional antiferromagnets, where magnetization and spin density vanish, the altermagnetic spin structure of RuO2 would yield distinct spin densities associated with opposite spin sublattices. That is, real-space spin densities of sublattices connected by the four-fold rotation symmetry around the [001] axis would differ [1]. However, in non-magnetic RuO2 bulk, the absence of local magnetic moments suggests that the antiferromagnetic order within each Ru plane of the (110) slab and the altermagnetic spin structure are unlikely. This is evident in the spin density visualization for the (110) slab displayed in Fig. 2(a) and the layer dependence of local moments on each Ru site from the surface to the bulk-like region shown in Fig. 2(c) (top and middle panels). Most atoms in the slab are non-magnetic except for those in the outermost layers.
In contrast to the bulk, local moments develop on the surface layers of the slab, decaying rapidly into the bulk. The collinear moments arrange ferrimagnetically within the surface layers, with unequal spin moments: Ru-center ($\sim$ 0.61 $\mu_{\mathrm{B}}$) and Ru-corner ($\sim$ $-0.17$ $\mu_{\mathrm{B}}$) couple collinearly in opposite directions. The majority spin density is located primarily on the Ru-center atom (denoted Ru-6f due to its six-fold coordination) and bridging oxygen (O-br, spin moment $\sim$ 0.19 $\mu_{\mathrm{B}}$), while the Ru-corner atom (denoted Ru-5f due to its five-fold coordination) dominates the minority spin density counterpart [cf. Fig. 2(b) for detailed labeling of atomic surface atoms]. Other oxygen ligands of the outermost layers exhibit non-spin-polarized behavior [Fig. 2(a)].
@@
As seen in Fig. 2(a) and evidenced by variations in the local moments for each Ru sublattice across layers, the magnetization density decreases dramatically from the surface layer to the innermost (bulk-like) layer. Local moments on Ru, and consequently the layer magnetization, nearly vanish starting from the layer just beneath the surface [Fig. 2(c), bottom panel]. The emergence of surface-layer magnetism and the rapid decrease in magnetization can also be observed in the layer-projected DOS in Fig. 2(e), where significant spin splitting occurs only in the top layer. This magnetic structure is more stable than the non-magnetic counterpart by approximately 45 meV per unit cell, independent of the slab thickness, thus emphasizing the origin of the magnetic structure due solely to the surface layer (see Fig. S2 in SM for the relative energy of the spin-polarized state compared to the non-spin-polarized counterpart as a function of slab thickness).
@@
In a recent study, Weber et al. [50] demonstrated that surface magnetization can spontaneously develop at the surface of an antiferromagnet with suitable symmetry, including the (110) surface of an AFM with a rutile structure, such as FeF2, which is isostructural to RuO2. Accordingly, surface magnetization is intrinsically linked to the magnetic structure of the bulk. However, as shown above, surface magnetization can arise solely from the local breaking of bulk symmetry at the surface of a non-magnetic material. Thus, the presence of surface magnetization alone does not constitute necessary or sufficient conditions to infer the bulk magnetic structure. Bulk measurements are required to identify the magnetic domain conclusively.
III.2.2 The origin of surface magnetization in RuO2 (110) surface
@@
In this section, we investigate the origin of surface magnetization on RuO2 (110) surface, which arises from the non-magnetic bulk state. When bulk RuO2 is transitioned to the (110) surface, the symmetry of the already distorted octahedral around the Ru-corner and Ru-center atoms of the surface layers is further reduced due to surface termination. The Ru-corner (Ru-5f) loses one coordination with one of its axially coordinated oxygen ligands, while the Ru-center (Ru-6f) maintains its coordination number. However, the two bridging oxygen ligands on the surface, which coordinate equatorially to the Ru-6f atom, become covalently unsaturated with only two bonds remaining as compared to three bonds in the bulk, contributing to structural distortions that affect the hybridization between Ru $4d$ and O $2p$ states. These distortions are expected to lift the degeneracy of the Ru $4d$ states further. Without any surface relaxation, i.e., no structural relaxation performed for slabs, the local symmetry of the Ru-5f is $C_{4v}$ (square pyramidal), while the Ru-6f retains its distorted octahedral symmetry as in the bulk. Nevertheless, this configuration is unstable, leading to structural relaxation akin to a Jahn-Teller distortion, which further lowers the local symmetry around the surface Ru atoms. During this relaxation, the Ru–O axial bonds in the Ru-5f-centered square pyramidal shorten, whereas the axial bonds in the Ru-6f-centered octahedra lengthen. Additionally, the equatorial Ru–O bonds centering around the Ru-6f become asymmetrically distorted; those involving the bridging oxygen ligands are significantly shortened, while the others elongate slightly. Despite these structural reconstructions, the total energy of the system remains high, with a considerable DOS at the Fermi level contributing to the occupation of anti-bonding states, as illustrated in Fig. 3(a).
@@
Figure 3: Chemical bonding analysis for the emergence of surface magnetization. (a) Comparison of PDOS around Fermi level in non-magnetic state, and spin-polarized state of the surface, (b) and (c) --pCOHP curves for Ru-O bonds around the Ru-6f (Ru-center) and Ru-corner (Ru-5f) atoms at the surface, respectively.
@@
To achieve stability, spin polarization is necessary to redistribute the two spin sublattices, resulting in spontaneous surface magnetization as shown in previous sections. Crystal orbital Hamilton population (COHP) analysis offers valuable insight into the driving forces behind the emergence of magnetization from a chemical bonding perspective [43, 44]. Fig. 3 presents the projected density of states for the atoms of the top surface, along with --pCOHP curves (averaged for the nearest-neighbor Ru–O contacts) for both structurally relaxed non-magnetic and magnetic states. In the non-magnetic state, RuO2 (110) displays a peak at the Fermi level in both the PDOS and --pCOHP (with negative values) curves, indicating an unstable configuration. This suggests that the non-magnetic surface might typically undergo structural relaxations, such as Peierls or Jahn-Teller instabilities. However, when we compared the relaxed geometrical structure of the non-spin-polarized surface with that of the spin-polarized counterpart, we observed no discernible structural difference. Instead, the surface exhibited pronounced electronic reorganization upon spin-polarization calculation. The surface atoms become magnetized and the electronic structure shows substantial spin-polarization [Fig. 2(e) and  3(a)]. Remarkably, this spin-polarized reorganization shifted the DOS and --pCOHP peaks away from the Fermi level, leading to a drastic decrease in the density of antibonding states at the Fermi level for both spin channels. Ultimately, it is this surface spin-polarization that stabilizes the entire structure.
@@
The development of surface magnetization on RuO2 (110) can also be understood through the lens of the Stoner criterion, which states that magnetization emerges if the product of the density of states at the Fermi level and the exchange interaction parameter exceeds unity [51]. In the non-magnetic state, the high DOS at the Fermi level suggests instability, driving the surface to electronic redistribution. Instead of structural relaxations like Peierls or Jahn-Teller distortions, the system minimizes its energy through spin polarization. This reorganization substantially lowers the DOS for both spin channels around the Fermi level as described above, stabilizing the surface. The significant reduction in antibonding states and the redistribution of electronic density through spin polarization effectively meets the Stoner criterion, leading to spontaneous surface magnetization.
@@
Electronic reconstruction that results in the emergence of induced magnetic moments and net magnetization in the surface layer can be further seen in the variation of charge states of individual atoms at the surface compared to their bulk counterparts, as detailed in Table S2. Different population methods yield varying values for these charge states, but the trend of electronic reconstruction for the surface atoms remains consistent across these methods. For example, using the charge density-based Bader approach, we find that the charge state of the Ru-5f atom decreases, while that of the Ru-6f atom increases relative to the bulk values. This is primarily due to the absence of one oxygen atom coordination of the Ru-5f (the axial oxygen), which leads to an increased electron density and a corresponding decrease in charge state at the Ru-5f site. In contrast, the Ru-6f retains full coordination with six oxygen ligands, but also specially interacts with two bridging oxygen ligands having higher hole densities due to one unsaturated coordination for each oxygen. This condition facilitates charge transfer from the Ru-6f to these covalently unsaturated ligands, thereby increasing its charge state.
III.2.3 The significance of surface magnetism in RuO2 (110) surface
@@
Surface magnetism developed on the surface of RuO2 (110) suggests that we can probe this phenomenon directly, i.e., the spin-splitting effect associated with surface bands should be detectable by surface- and spin-sensitive spectroscopy techniques such as spin-resolved ARPES. Fig. 4 shows the slab band structure with weighted contribution from the top layer atoms (2 Ru and 4 O, cf. Fig. 2(d)) along with projected bulk bands in the light-gray colored background. Interestingly, low energy electronic structures close to the Fermi level exhibit surface states with flatness and significant spin-splitting exchange energy of approximately 0.60 eV, as indicated by black arrows in Fig. 4(a) and  4(b). The occupied flat bands, which lie  0.40 eV below Fermi level, are derived from all magnetic atoms of the surface layer, i.e., Ru-5f, Ru-6f, and O-br, with the strongest contribution from Ru-6f followed by O-br and the lowest contribution from Ru-5f (cf. Fig. S3 in SM). Other non-magnetic atoms of the top layer such as O-3f and O-sub contribute negligibly to these states. The minority spin counterparts of these flat bands are unoccupied (located  0.20 eV above Fermi level) and dominated by the Ru-6f character and a smaller contribution from O-br (cf. Fig. S3 in SM). This is consistent with the highest calculated magnetic moment value for Ru-6f atom amongst others on the surface.
@@
Figure 4: Projected electronic band structure for the atoms of the outermost layer plotted along the corresponding projected bulk bands shown in the light-gray background (Ef set to 0). Each layer contains 2 formula units of RuO2. (a) Majority spin (spin up, red), (b) minority spin (spin down, blue), (c) non-spin-polarized band. Surface bands can be seen with high intensity of weighted data, those near the Fermi level pointed to by black arrows.
@@
Interestingly, when studying the origin of these surface states along both the [$01\overline{1}$ 0] direction and the perpendicular in-plane [001] direction, and conditions under which these states exist (dubbed flat band surface state—FBSS), Jovic et al. [52] assigned the existence of FBSS due mainly to bridging oxygen atoms. Consequently, under Ru-rich condition, these atoms are removed and the surface states are expected to disappear. However, our result shows that these surface states are derived from both the Ru and the bridging oxygen atoms of the outermost layers, providing insight into the origin of the flat bands near the Fermi level. It should be noted that the presence of occupied flat bands near the Fermi level has been observed in previous experiments  [8, 23, 53, 54]. However, the spin polarization of these bands and the presence of their spin exchange counterpart above the Fermi level have not been reported in the literature. This is likely due to the ARPES technique probing only occupied states. Therefore, unoccupied states-sensitive experiments such as inverse ARPES, scanning tunneling microscopy (STM), or X-ray absorption spectroscopy (XAS) might be required to realize the full spin-polarized spectrum of the flat bands in the proximity of the Fermi level due to the surface magnetism in RuO2 (110).
@@
As mentioned above, the existence of flat bands lying closely beneath the Fermi level has been observed experimentally, but the origin of these states is still a debate in the literature. Jovic et al. [54] showed that the flat bands have a surface origin. They were considered to be topologically trivial states that connect the projection of nodal lines along the XR directions in the bulk BZ [54]. Very recently, based on spin-resolved ARPES data, people also assigned the flat bands near the Fermi level along the $\Gamma\mathrm{M}$ momentum direction as surface states [10]. However, a recent experiment has also shown the existence of flat bands along that $\Gamma\mathrm{M}$ direction, but the authors argued that it likely has a bulk origin [23]. It is noteworthy that our projected bulk spectrum exhibits a handful of flat bands residing below the Fermi level and overlapping in energy with the flat bands of interest, i.e., those with significant contribution from the outermost layers as described above. This overlap could be the source of confusion in the literature. To reconcile the origin of these flat bands, we calculated the layer-resolved (projected) band structure for each layer of the slab, and the results were shown in Fig. S4 of SM. The layer-projected band structures show a rapid decrease in the contribution of each layer to the flat bands of interest when going from the outermost to the innermost (or the middle, the bulk-like) layer. The middle layer of the slab displays no fingerprint of the pairs of bands being discussed (Fig. S4(f)), suggesting that these bands are derived from the top and bottom layers (the surface layers). In addition, the charge density associated with the flat bands at the $\overline{\Gamma}$ point plotted in Fig. S5 of SM shows a clear outermost-layers-origin of the bands, further corroborating the surface nature of them. The fact that those surface bands are spin-split is clear evidence for the emergence of surface magnetization mostly due to the atoms at the surface. This is consistent with the observation that the developed local moments on atoms decay quickly when going from surface to the bulk region.
@@
Another pair of spin-splitting bands can be observed around --1.5 eV where the minority spin channel is completely flat while the majority counterpart is slightly dispersive. This contrasts with the corresponding surface-dominated bands of the non-magnetic surface where these bands are also slightly dispersive in the proximity of the $\overline{\Gamma}$ point but without a spin polarization (Fig. 4(c)). Similar to the flat band surface states close to the Fermi level, the spin-polarized states split in energy, i.e., one moving up and the other moving down compared to non-magnetic states upon spin-polarization calculation. In addition, there is another set of surface states sitting at  1.5 eV above the Fermi level. These states are slightly dispersive near the $\overline{\Gamma}$ point and do not show a significant exchange splitting, likely due to the dominating contribution of the small moment Ru-5f atoms to the band character.
@@
Considering that the bulk RuO2 is non-magnetic and our calculations exclude relativistic effects, the emergence of significant spin exchange splitting associated with surface bands near the Fermi level underscores the unique feature of the spontaneous surface magnetization of RuO2 (110). This characteristic is likely a key factor in determining the magnetic properties of surface RuO2. Additionally, the presence of large spin-splitting surface bands, particularly those close to the Fermi level, suggests that the spin polarization of transport at the surface would differ significantly from that in the bulk, which could profoundly influence spin-dependent phenomena in RuO2. Therefore, we conjecture that surface magnetism can play a vital role in the material’s spin Hall magnetoresistance when coupled with heavy metals such as Pt or Ta, due to the induced exchange interactions and potential spin current manipulation at the surface.

### v2

@@ III.2.1 Emergence of surface magnetization
Figure 2: Geometrical structure of the (110) surface and the emergence of surface magnetism. (a) The slab structure of 13L RuO2(110) includes two Ru sublattices inherited from the bulk: Ru-center in red, Ru-corner in blue, and oxygen ligands in silver. At the surface layers, Ru-center atoms are fully octahedrally coordinated by six O ligands (denoted as Ru-6f), while the Ru-corner sublattice has one dangling bond due to surface termination (Ru-5f), see (b) for a better view. Also shown is the magnetization density visualization, which exhibits significant accumulation of spin moments only at the outermost layers, with spin-up in yellow and spin-down in cyan. The isosurfaces are plotted using 5% of the maximum isosurface value. (b) Detailed atomic structure of the surface includes stoichiometric surface termination, with labels for atoms in the surface layer. (c) The spin moment at each Ru site in each layer (top two panels), and the magnetization of each layer (bottom panel) are plotted as functions of distance from the bulk-like region. (d) View along the [001] direction showing the layers numbered for the layer-projected DOS in (e), demonstrating that spin polarization decays quickly when approaching the inner layers.
@@
To gain insight into the magnetic structure of the RuO2(110) surface, we performed calculations using the slab model with 3D periodic boundary condition. We simulate a surface with a bulk-like region well represented in the middle of the structure, choosing stoichiometric slabs with an odd number of Ru layers. These slabs possess inversion symmetry, eliminating unnecessary complications related to asymmetric systems, such as spurious charge transfer and build up electric fields across the slab thickness. The slab thickness ranges from three layers (denoted as 3L) up to nineteen layers (19L). Each layer is composed of three atomic planes, with one Ru-containing layer sandwiched between two oxygen-only layers. Each slab has two identical surface layers at the top and bottom of the unit cell [cf. Fig. 2(a)].
@@
The converged, relaxed thirteen-layer slab unit cell (13L) is shown in Fig. 2(a). This symmetric structure represents both bulk and surface regions since it is thick enough to decouple the two surfaces, and the middle layer behaves as a bulk-like region [see Fig. S1 in the Supplemental Material (SM) for the dependence of local moments of surface Ru atoms and top-layer magnetization on slab thickness]. Each layer stacked along the [110] direction contains two Ru atoms corresponding to the Ru-center and Ru-corner sublattices of the bulk [cf. Fig. 1(a)]. This arrangement would yield a magnetically compensated structure with oppositely pointed moments within each (110) plane if the ground state of the bulk was antiferromagnetically ordered with the Néel vector along [001], as assumed earlier—similar to the case of the isostructural FeF2 antiferromagnet [54].
@@
Unlike the magnetic structure in traditional antiferromagnets, where magnetization and spin density vanish, the assumed altermagnetic spin structure of RuO2 would yield distinct spin densities associated with opposite spin sublattices. That is, real-space spin densities of sublattices connected by the four-fold rotation symmetry around the [001] axis would differ [1]. However, in non-magnetic RuO2 bulk, the absence of local magnetic moments suggests that the antiferromagnetic order within each Ru plane of the (110) slab and the altermagnetic spin structure are unlikely. This is evident in the spin density visualization for the (110) slab displayed in Fig. 2(a) and the layer dependence of local moments on each Ru site from the surface to the bulk-like region shown in Fig. 2(c) (top and middle panels). Most atoms in the slab are non-magnetic except for those in the outermost layers.
[…]
@@
As seen in Fig. 2(a) and evidenced by variations in the local moments for each Ru sublattice across layers, the magnetization density decreases dramatically from the surface layers to the bulk-like layers. Local moments on Ru, and consequently layer magnetizations, nearly vanish starting from the layer just beneath the surface [Fig. 2(c), bottom panel]. The emergence of surface-layer magnetism and the rapid decrease in magnetization can also be observed in the layer-projected DOS in Fig. 2(e), where significant spin splitting is shown to occur mainly on the top layer. This magnetic structure is lower in energy than the non-magnetic counterpart by $\sim$45 meV per unit cell, independent of the slab thickness, consistent with the origin of the magnetic structure being solely due to the surface layer (see Fig. S2 for the relative energy of the spin-polarized state compared to the non-spin-polarized counterpart as a function of slab thickness).
@@
In a recent study, Weber et al. [54] demonstrated that surface magnetization can spontaneously emerge at the surface of an antiferromagnet with suitable symmetry, including the (110) surface of an AFM with a rutile structure, such as FeF2, which is isostructural to RuO2. Accordingly, surface magnetization is intrinsically linked to the magnetic structure of the bulk. However, as shown above, surface magnetization also arises from the local breaking of bulk symmetry at the surface of a non-magnetic material. Further, we want to emphasize that the surface magnetism seen in our simulation is robust with DFT functionals as long as the functional gives rise to non-magnetic bulk state, i.e., those PBE+$U$ with the on-site $U$ values not exceeding 1.06 eV [6]. Thus, the presence of surface magnetization alone does not constitute necessary or sufficient conditions to infer the bulk magnetic structure. Bulk measurements are required to identify the magnetic domain conclusively.
III.2.2 The cause of surface magnetization in RuO2(110)
@@
In this section, we investigate the origin of surface magnetization on RuO2(110), considering the non-magnetic bulk state. From bulk RuO2 to the (110) surface, the symmetry of the already distorted octahedral around the Ru-corner and Ru-center atoms of the surface layers is further reduced due to surface termination. The Ru-corner (Ru-5f) loses one of its axially coordinated oxygen ligands, while the Ru-center (Ru-6f) maintains its coordination number. The two bridging oxygen ligands on the surface, which coordinate equatorially to the Ru-6f atom, become covalently unsaturated with only two bonds remaining as compared to three bonds in the bulk, contributing to structural distortions that affect the hybridization between Ru $4d$ and O $2p$ states. These distortions are expected to lift the degeneracy of the Ru $4d$ states further. Without any surface relaxation, i.e., no structural relaxation performed for slabs, the local symmetry of the Ru-5f is $C_{4v}$ (square pyramidal), while the Ru-6f retains its distorted octahedral symmetry as in the bulk. Nevertheless, this configuration is unstable, leading to structural relaxation akin to a Jahn-Teller distortion, which further lowers the local symmetry around the surface Ru atoms. During this relaxation, the Ru–O axial bonds in the Ru-5f-centered square pyramidal shorten, whereas the axial bonds in the Ru-6f-centered octahedra lengthen. Additionally, the equatorial Ru–O bonds centering around the Ru-6f become asymmetrically distorted; those involving the bridging oxygen ligands are significantly shortened, while the others elongate slightly. Despite these structural reconstructions, the total energy of the system remains high, with a considerable DOS at the Fermi level contributing to the occupation of anti-bonding states, as illustrated in Fig. 3(a).
@@
Figure 3: Chemical bonding analysis for the emergence of surface magnetization. (a) Comparison of PDOS around Fermi level in the non-magnetic state, and spin-polarized state of the surface, (b) --pCOHP curves averaged for Ru–O bonds around the Ru-6f (Ru-center), and (c) around the Ru-5f (Ru-corner) atoms at the surface ($E_{\text{F}}$ set to 0).
@@
To achieve stability, spin polarization is necessary to redistribute electronic densities of the two spin sublattices, resulting in spontaneous surface magnetization as shown in the previous section. Crystal orbital Hamilton population (COHP) analysis offers valuable insight into the driving forces behind the emergence of magnetization from a chemical bonding perspective [47, 48]. Fig. 3 presents the projected density of states (PDOS) on the atoms of the top surface, along with --pCOHP curves (averaged for the nearest-neighbor Ru–O contacts) for both structurally relaxed non-magnetic and magnetic states. In the non-magnetic state, despite the structural relaxation, RuO2(110) displays a peak at the Fermi level in both the PDOS and --pCOHP (with negative values) curves, indicating an unstable configuration. This suggests that structural effect alone is not enough to stabilize the surface system. Consequently, the surface exhibited pronounced electronic reorganization upon spin-polarization calculation. The surface atoms become magnetized and the electronic structure shows substantial spin-polarization [Fig. 2(e) and  3(a)]. Remarkably, this spin-polarized reorganization shifted the DOS and --pCOHP peaks away from the Fermi level, leading to a drastic decrease in the density of antibonding states at the Fermi level for both spin channels, and effectively stabilizing the system by about 43 meV. The electronic redistribution effect is stronger for Ru-6f compared to that of Ru-5f, thus giving rise to larger moments on the former as seen in the previous section. In addition, when performing further structural relaxation with the inclusion of spin-polarization, we observed no discernible structural difference to the non-spin-polarized case, gaining only 2 meV in total energy. Ultimately, it is the surface spin-polarization that stabilizes the surface by about 45 meV per unit cell, regardless of the slab thickness, as mentioned in the previous section.
@@
The development of surface magnetization on RuO2(110) can also be understood through the lens of the Stoner criterion, which states that magnetization emerges if the product of the density of states at $E_{\text{F}}$ and the exchange interaction parameter exceeds unity [55]. In the non-magnetic state, even after the structural relaxation, the high DOS at the Fermi level suggests the surface instability. This unstable state drives the surface to electronic redistribution, and the system minimizes its energy through spin polarization. This reorganization substantially lowers the DOS for both spin channels around $E_{\text{F}}$ as described above, stabilizing the surface. The significant reduction in antibonding states and the redistribution of electronic density through spin polarization effectively meet the Stoner criterion, leading to spontaneous surface magnetization.
@@
Electronic reconstruction that results in the emergence of induced magnetic moments and net magnetization in the surface layer can be further seen in the variation of charge states of individual atoms at the surface compared to their bulk counterparts, as detailed in Table S2. Different population methods yield varying values for these charge states, but the trend of electronic reconstruction for the surface atoms remains consistent across these methods. For example, using the charge density-based Bader approach, we find that the charge state of the Ru-5f atom generally decreases, while that of the Ru-6f atom increases relative to the bulk values. This is primarily due to the absence of one oxygen atom coordination (the axial oxygen) of the Ru-5f, which leads to an increased electron density and a corresponding decrease in charge state at the Ru-5f site. In contrast, the Ru-6f retains full coordination with six oxygen ligands, but also specially interacts with two bridging oxygen ligands having higher hole densities due to one unsaturated coordination for each oxygen. This condition facilitates charge transfer from the Ru-6f to these covalently unsaturated ligands, thereby increasing its charge state.
III.2.3 The significance of surface magnetism in RuO2(110)
@@
Surface magnetism developed on the surface of RuO2(110) suggests that we can probe this phenomenon directly, i.e., the spin-splitting effect associated with surface bands should be detectable by surface- and spin-sensitive spectroscopy techniques such as spin-resolved ARPES. Fig. 4 shows the slab band structure with weighted contribution from the top layer atoms (2 Ru and 4 O, cf. Fig. 2(d)) along with projected bulk bands in the light-gray colored background. Interestingly, low energy electronic structures close to the Fermi level exhibit surface states with flatness and significant spin-splitting exchange energy of approximately 0.60 eV, as indicated by black arrows in Fig. 4(a) and  4(b). The occupied flat bands, which lie approximately 0.40 eV below $E_{\text{F}}$, are derived from all magnetic atoms of the surface layer, i.e., Ru-5f, Ru-6f, and O-br, with the strongest contribution from Ru-6f followed by O-br and the lowest contribution from Ru-5f, which is consistent with previous observation [56] (cf. Fig. S3 in SM). Other non-magnetic atoms of the top layer such as O-3f and O-sub contribute negligibly to these states. The minority spin counterparts of these flat bands are unoccupied (located  0.20 eV above $E_{\text{F}}$) and dominated by the Ru-6f character and a smaller contribution from O-br (cf. Fig. S3 in SM). This is consistent with the highest calculated magnetic moment value for the Ru-6f atom among others on the surface.
@@
Figure 4: Projected electronic band structure for the atoms of the outermost layer plotted along the corresponding projected bulk bands shown in the light-gray background ($E_{\text{F}}$ set to 0). Each layer contains 2 formula units of RuO2. (a) Majority spin (spin up, red), (b) minority spin (spin down, blue), (c) non-spin-polarized band. Surface bands can be seen with high intensity of weighted data, those near the Fermi level pointed to by black arrows.
@@
It should be noted that the presence of occupied flat bands near the Fermi level has been observed in previous experiments by using ARPES [10, 26, 57, 58]. However, the spin polarization of these bands and the presence of their spin exchange counterpart above $E_{\text{F}}$ have not been reported in the literature. This is likely due to the ARPES technique probing only occupied states. Therefore, unoccupied states-sensitive experiments such as inverse ARPES, scanning tunneling microscopy (STM), or X-ray absorption spectroscopy (XAS) might be required to realize the full spin-polarized spectrum of the flat bands in the proximity of the Fermi level due to the surface magnetism in RuO2(110).
@@
As mentioned above, the existence of flat bands lying closely beneath the Fermi level has been observed experimentally, but the origin of these states is still a debate in the literature. Jovic et al. [58] showed that the flat bands have a surface origin. They were considered to be topologically trivial states that connect the projection of nodal lines along the XR directions in the bulk BZ [58]. Very recently, based on spin-resolved ARPES data, Liu et al. [10] assigned the flat bands near the Fermi level along the $\Gamma$$\mathrm{M}$ momentum direction as surface states. However, a recent experiment has also shown the existence of flat bands along that $\Gamma$$\mathrm{M}$ direction, but the authors argued that it likely has a bulk origin [26]. It is noteworthy that our projected bulk spectrum exhibits a few flat bands residing below the Fermi level and overlapping in energy with the flat bands of interest, i.e., those with significant contributions from the outermost layers as described above. This overlap could be the source of confusion in the literature.
To reconcile the origin of these flat bands, we calculated the layer-resolved (projected) band structure for each layer of the slab, and the results are shown in Fig. S4 of SM. The layer-projected band structures show a rapid decrease in the contribution of each layer to the flat bands of interest when going from the outermost to the innermost (or the middle, the bulk-like) layer. The middle layer of the slab displays no fingerprint of the pairs of bands being discussed (Fig. S4(f)), suggesting that these bands are derived from the top and bottom layers (the surface layers). In addition, the charge density associated with the flat bands at the $\overline{\Gamma}$ point plotted in Fig. S5 clearly shows the outermost-layer origin of the bands. The fact that those surface bands are spin-split is a clear evidence for the emergence of surface magnetization mostly due to the atoms at the surface. This is consistent with the observation that the developed local moments on atoms decay quickly when going from the surface to the bulk region.
@@
Another pair of spin-splitting bands can be observed around --1.5 eV where the minority spin channel is completely flat while the majority counterpart is slightly dispersive. This contrasts with the corresponding surface-dominated bands of the non-magnetic surface where these bands are also slightly dispersive in the proximity of the $\overline{\Gamma}$ point but without spin polarization (Fig. 4(c)). Similar to the flat band surface states close to the Fermi level, these spin-polarized states split in energy, i.e., one moving up and the other moving down compared to non-magnetic states upon spin-polarization calculation. In addition, there is another set of surface states sitting at around 1.5 eV above the Fermi level. These states are slightly dispersive near the $\overline{\Gamma}$ point and do not show a significant exchange splitting, likely due to the dominating contribution of the small moment Ru-5f atoms to the band composition.
@@
Considering that the bulk RuO2 is non-magnetic and our calculations exclude relativistic effects, the emergence of significant spin exchange splitting associated with surface bands near the Fermi level underscores the unique feature of the spontaneous surface magnetization of RuO2(110). This characteristic is likely a key factor in determining the magnetic properties of surface RuO2. Additionally, the presence of large spin-splitting surface bands, particularly those close to the Fermi level, suggests that the spin polarization of transport at the surface would differ significantly from that in the bulk, which could profoundly influence spin-dependent phenomena in RuO2. Therefore, due to the induced exchange interactions, proximity effect, and potential spin current manipulation at the surface, we conjecture that surface magnetism could play a vital role in the material’s spin Hall magnetoresistance when coupled with heavy metals such as Pt or Ta [27], or serve a relevant source for interfacial exchange coupling when paired with a magnetic material.